EPF-Lausanne/15 July 2008

From 2008.igem.org

<<Previous - Back to Notebook - Next>>

The transformations of yesterday didn't work, except two. There where fast no bacteria on our plates, the control also didn't work. We recovered 4 colonies each from the Prey1 in CM (and now we know which is which), and PLD plasmid in Kanamycin. We grow them overnight shaking at 37°C. The rest we left downstairs at 37°C on the plates, hoping perhaps they are just slow growers.

We have done some glycerol stocks with the predator1 and predator2. (the one we let grow overnight from yesterday). Stock was also done for Top10 cells mistakenly grown in LB (emergency stock) and for Top10 cells grown in SOB for competence. We had 4 samples of Top10, we grew one divided into 2 mL each mL in 250 mL of SOB (pH 7.6). One of the samples was put in shaker, the other one was left outside alone (if this works it will free space in the incubator of the lab, which we share with others who sometimes need it). Incubation started at 17h45, and should be ready to use 16 hours later, at 9h45 tomorrow.

We try to do a new transformation :

- Cells : DH5alpha
- Plasmids : pHD1 (from Ron Weiss system) with CM R, rhlR (part BBa_I0466 : 5c 1001) 

with Amp R and rhlI (part BBa_C0070 : 3E 1014) (from iGEM parts) with Kan R, and pUC19 for the control with Amp R.

We need some new SOB.